OpenStudy (anonymous):

3-4(z+1)=3z-(2+z) Solve for Z

OpenStudy (anonymous):

Here's how you could solve this. First, expand the brackets \[3 - 4 z - 4 = 3 z -2 -z\] Which you simplify to: \[ -4 z -1 = 2z -2 \] Now add \(4 z\) to both sides of the equation: \[ -1 = 6 z -2 \] And add 2 to both sides: \[ 1 = 6 z \] Divide by \( 6 \) \[ 1/6 = z \] So \( z = 1/6 \). Insert this into the original equation to check.

OpenStudy (anonymous):

3-4(z+1)=3z-(2+z) 3-4z+4=3z-2-z z=1/6 check; 3-4(1/6+1)=3(1/6)-(2+1/6) -5/3=-5/3